OpenStudy (anonymous):

Brad wants to buy flowers for his friend with $45...? ....The daises are $1 each and the roses are $2 each. He buys 3 more daises than roses. How much did the roses cost? Notice obviousness that the answer is not 2$, need whole problem.

OpenStudy (jackellyn):

d=r+3 45=2r+d Substitute for d in second equation 45=2r+(r+3) 45=2r + r + 3 45= 3r + 3 45-3=3r 42=3r 42/3=r 14=r d=(14)+3 d=17 Check: 45=2(14)+17 45=28+17 45=45 Roses are 2r 2r=2(14)=28

OpenStudy (jackellyn):

@chrisfig17